# DeepXiv-SDK: An Agentic Data Interface for Scientific Papers

### v1

@@ Abstract
Research agents are increasingly used in AI4Science for scientific information seeking and evidence-grounded decision making. Yet a persistent bottleneck is paper access: agents typically retrieve PDF/HTML pages, heuristically parse them, and ingest long unstructured text, leading to token-heavy reading and brittle evidence lookup. This motivates an agentic data interface for scientific papers that standardizes access, exposes budget-aware views, and treats grounding as a first-class operation.
We introduce DeepXiv-SDK, which enables progressive access aligned with how agents allocate attention and reading budget. DeepXiv-SDK exposes as structured views a header-first view for screening, a section-structured view for targeted navigation, and on-demand evidence-level access for verification. Each layer is augmented with enriched attributes and explicit budget hints, so agents can balance relevance, cost, and grounding before escalating to full-text processing. DeepXiv-SDK also supports multi-faceted retrieval and aggregation over paper attributes, enabling constraint-driven search and curation over paper sets.
DeepXiv-SDK is currently deployed at arXiv scale with daily synchronization to new releases and is designed to extend to other open-access corpora (e.g., PubMed Central, bioRxiv). We release RESTful APIs, an open-source Python SDK, and a web demo showcasing deep search and deep research workflows; the service is free to use with registration.
DeepXiv-SDK: An Agentic Data Interface for Scientific Papers
Hongjin Qian, Ziyi Xia, Ze Liu, Jianlv Chen, Kun Luo, Minghao Qin Chaofan Li, Lei Xiong, Sen Wang, Zhengyang Liang, Zheng Liuthanks: Corresponding author. Beijing Academy of Artificial Intelligence {chienqhj,zhengliu1026}@gmail.com Project Page: https://github.com/DeepXiv/deepxiv_sdk
@@ 2.2 Data Layer: Corpus-Scale Ingestion, Structuring, and Signal Materialization
Capability Endpoint (template) Returns (role)
Progressive access (arXiv)
Quick Brief /arxiv (type=brief, id={id}) Brief metadata.
Retrieval and utilities
Corpus extensibility (PMC)
@@ 3 License
DeepXiv-SDK is designed as an agentic access and structuring interface rather than a full-text redistribution service. We reuse arXiv descriptive metadata (e.g., title, authors, abstract, categories, and timestamps) under the arXiv API Terms of Use, which state that arXiv e-print metadata are available under CC0 1.0.111https://info.arxiv.org/help/api/tou.html Beyond metadata, DeepXiv-SDK derives navigation and enrichment signals (e.g., section structure, summaries, budget cues, and optional scholarly/social context) to support agentic screening, routing, and evidence-seeking.
@@ Appendix A Implementation and Deployment
DeepXiv-SDK consists of an arXiv-scale processing pipeline that normalizes papers into section-addressable objects with enriched signals, and a serving stack that exposes progressive access and hybrid retrieval through a cached REST interface. The motivation is to replace repeated per-agent PDF parsing with a once-for-all curation layer that is stable, budget-aware, and reusable across agent workflows.
@@ Appendix B Statistics
Question Answer (arXiv ID)
@@
14 "Peking University",
15 "Beijing Academy of Artificial Intelligence",
16 "Beijing, China"
17 ]
18 }...
19 ],
@@
28 "tldr": "MemoRAG introduces a novel dual-system framework that mimics human cognition by first building a global memory of long documents and then using memory-derived retrieval clues to efficiently and accurately locate relevant information, solving the limitations of traditional RAG in handling unstructured, long-context tasks without explicit search intent.",
29 "token_count": 1661
30 }...
31 ],
32 "categories": [
33 "cs.CL",
34 "cs.AI"
35 ],
@@
39 "long-context retrieval",
40 "reinforcement learning from generation feedback",
41 "KV compression in LLMs",
42 "unstructured document processing"
@@
44 "tldr": "[research paper] MemoRAG introduces a dual-system RAG framework that mimics human cognition by first constructing a global memory of long contexts via a lightweight...",

### v2

# arXiv:2603.00084v2[cs.DL] 03 Mar 2026

@@ Abstract
LLM-agents are increasingly used to accelerate the progress of scientific research. Yet a persistent bottleneck is data access: agents not only lack readily available tools for retrieval, but also have to work with unstrcutured, human-centric data on the Internet, such as HTML web-pages and PDF files, leading to excessive token consumption, limit working efficiency, and brittle evidence look-up. This gap motivates the development of an agentic data interface, which is designed to enable agents to access and utilize scientific literature in a more effective, efficient, and cost-aware manner.
In this paper, we introduce DeepXiv-SDK, which offers a three-layer agentic data interface for scientific literature. 1) Data Layer, which transforms unstructured, human-centric data into normalized and structured representations in JSON format, improving data usability and enabling progressive accessibility of the data. 2) Service Layer, which presents readily available tools for data access and ad-hoc retrieval. It also enables a rich form of agent usage, including CLI, MCP, and Python SDK. 3) Application Layer, which creates a built-in agent, packaging basic tools from the service layer to support complex data access demands.
DeepXiv-SDK currently supports the complete ArXiv corpus, and is synchronized daily to incorporate new releases. It is designed to extend to all common open-access corpora, such as PubMed Central, bioRxiv, medRxiv, and chemRxiv. We release RESTful APIs, an open-source Python SDK, and a web demo showcasing deep search and deep research workflows. DeepXiv-SDK is free to use with registration.
DeepXiv-SDK: An Agentic Data Interface for Scientific Literature
Hongjin Qian, Ziyi Xia, Ze Liu, Jianlyu Chen, Kun Luo, Minghao Qin, Chaofan Li, Lei Xiong, Junwei Lan, Sen Wang, Zhengyang Liang, Yingxia Shao, Defu Lian, Zheng Liuthanks: Corresponding author. Beijing Academy of Artificial Intelligence {chienqhj,zhengliu1026}@gmail.com Project Page: https://github.com/DeepXiv/deepxiv_sdk
@@ 1 Introduction
2 Related Work
As LLM agents increasingly rely on tool use and multi-step interaction, data access often becomes a bottleneck for scaling agent capability, since agents must repeatedly retrieve, parse, and ground on external artifacts such as web pages and PDFs (Wang et al., 2024b; Yao et al., 2023; OpenAI, 2025). A growing line of work therefore argues for turning raw web artifacts into structured, agent-callable interfaces with normalized schemas, controllable views, and provenance to reduce ad hoc parsing and improve reliability (Qian and Liu, 2025; Song et al., 2025). Within scientific literature, many prior efforts focus on agent frameworks or web platforms to improve literature review efficiency (e.g., search-then-read workflows, iterative summarization, and multi-document QA), but typically do not expose a reusable data interface API that agents can call deterministically across tasks (He et al., 2025; Miao et al., 2025; Ju et al., 2025). The closest efforts in spirit are ar5iv and AlphaXiv, which provide more usable browsing experiences by converting arXiv papers into readable HTML and enriched views. [^1] https://ar5iv.labs.arxiv.org/ [^2] https://www.alphaxiv.org/ These systems are best viewed as advanced mirrors of arXiv for human browsing, rather than agentic data interfaces designed to scale through a reusable API and progressive, budget-aware access.
In contrast, DeepXiv-SDK provides an accessible, progressive interface with explicit budget cues and hybrid, attribute-conditioned retrieval, enabling agents to screen cheaply, read selectively, and verify on demand.
@@ 4 License
DeepXiv-SDK is designed as an agentic access and structuring interface rather than a full-text redistribution service. We reuse arXiv descriptive metadata (e.g., title, authors, abstract, categories, and timestamps) under the arXiv API Terms of Use, which state that arXiv e-print metadata are available under CC0 1.0.333https://info.arxiv.org/help/api/tou.html Beyond metadata, DeepXiv-SDK derives navigation and enrichment signals (e.g., section structure, summaries, budget cues, and optional scholarly/social context) to support agentic screening, routing, and evidence-seeking.
@@ 6 Conclusion
Implementation details of DeepXiv-SDK (data processing and service deployment) are provided in Appendix A, current corpus statistics are summarized in Appendix B, and Listing 1 showcases the metadata format returned for an example paper.
@@ Appendix A Implementation and Deployment
DeepXiv-SDK consists of an arXiv-scale processing pipeline that normalizes papers into section-addressable objects with enriched signals, and a serving stack that exposes progressive access and hybrid retrieval through a cached REST interface.
@@ Appendix B Statistics
Capability Endpoint (template) Returns (role)
Progressive access (arXiv)
Quick Brief /arxiv (type=brief, id={id}) Brief metadata.
Retrieval and utilities
Corpus extensibility (PMC)
@@
10 "Beijing Academy of Artificial Intelligence", ]}...],
@@
19 "tldr": "MemoRAG introduces a novel dual-system framework that mimics human cognition by first building a global memory of long documents...",
20 "token_count": 1661 }... ],
21 "categories": ["cs.CL", "cs.AI"],
@@
25 "long-context retrieval", ...
@@
27 "tldr": "[research paper] MemoRAG introduces a dual-system RAG framework that mimics human cognition by first constructing a global memory of long...",